# Is Stephen Curry really a guard? New perspective on players typology using functional data analysis

### v1

@@ 1 Introduction
With the increasing availability of large, detailed, and complex sports performance datasets, sports analytics has become a rapidly expanding field of both research and application, drawing significant interest from the statistics and machine learning communities. In team sports, traditional approaches have often focused on predicting game outcomes based on team characteristics (Dixon and Coles,, 1997; Baboota and Kaur,, 2019); for recent reviews, see Horvat and Job, (2020); Bunker and Susnjak, (2022). More recently, the availability of tracking data has enabled researchers to delve into more specific aspects of various sports, such as evaluating individual player performance in soccer (Pappalardo et al.,, 2019; Decroos et al.,, 2019) or analyzing defensive performance in American football passing plays (Schmid et al.,, 2021). See Kovalchik, (2023) for a recent review on machine learning approaches to analyze tracking data. Basketball, in particular, presents opportunities for analysis, supported by extensive data provided by the National Basketball Association (NBA).
@@
Basketball is a team sport in which two teams of five players compete to score points by shooting a ball through the opponent’s hoop while defending their own. The game has been extensively studied through various analytical perspectives. For example, researchers have examined shooting mechanics and movements (Miller and Bartlett,, 1996; Liu and Burton,, 1999; Okazaki et al.,, 2015), estimated pre-game and in-game winning probabilities (Maddox et al.,, 2022), and modeled outcomes of possessions (Cervone et al.,, 2016). Statistical models have also been applied to analyze defensive strategies (Csataljay et al.,, 2013) and spatial performance on the court (Zuccolotto et al.,, 2023). For a comprehensive review of performance analysis in basketball, see Terner and Franks, (2021).
@@
In basketball, game outcomes are heavily influenced by shooting accuracy, as a player’s effectiveness is often defined by its ability to consistently score from various court locations. Analyzing shooting patterns thus serves as a powerful tool, not only for assessing individual performance but also for refining team strategies. An important tool for examining shooting patterns is shooting charts, which are visual representations that display all the shots a player has attempted and made over a season or a single game. Shooting charts are extensively studied in basketball research, with much of the work focusing on modeling them as spatial point processes. This approach typically involves creating a count matrix of shots taken by each player on a discretized court, then estimating an intensity surface over this grid by fitting a log-Gaussian Cox process. A low-rank approximation of these intensity surfaces is achieved using non-negative matrix factorization, as described in Miller et al., (2014); Franks et al., (2015). Non-parametric estimation of the intensity surface is also considered in Yin et al., (2022). These low-rank intensity surfaces help characterize NBA players’ shooting tendencies and evaluate their scoring efficiency. Some models do not require the intensity surface fitting approach. For example, Reich et al., (2006) employed conditionally autoregressive models to capture spatial correlations, while Hu et al., (2023) fitted a Bayesian zero-inflated Poisson model to represent shooting attempts for players with distinct shooting styles. Although these methods effectively summarize player data, they are less suitable for comparing different players.
@@
To compare players’ shooting behaviors, specialized methods have been developed. Fu and Stasko, (2024) employ (interactive) data visualization techniques to analyze and contrast shooting performance. In a different approach, Hu et al., (2021) use a log-Gaussian Cox process to model the shooting intensity with a specific similarity measure for comparing players. Jiao et al., (2021) propose a Bayesian marked spatial point process model to capture shot intensity, investigating whether players’ success rates are higher in areas where they attempt more shots and clustering players based on this pattern. In another study, Wong-Toi et al., (2022) analyze shooting patterns across players by dividing the court into a grid and using a Bayesian nonparametric mixture model to identify latent clusters among players. Similarly, Yin et al., (2023) apply a Bayesian nonparametric matrix clustering technique to estimated intensity maps, offering insights into shot selection characteristics. All these methods require a discretized representation of the court.
@@
To overcome the requirement of court discretization, we propose to estimate the shot density continuously across the offensive half-court and model these densities as functional data (Ramsay and Silverman,, 2005; Kokoszka and Reimherr,, 2017). This idea first appeared in our paper Golovkine et al., (2024). Both missed and made shot charts are treated as multivariate functional data over a two-dimensional domain (i.e., the offensive half-court). Functional Data Analysis (FDA) addresses the analysis of realizations of a continuous stochastic process, observed at discrete sampling points and defined on compact domains. By treating shot density as a stochastic process defined over the offensive half-court, we can consider each player’s shot density as a realization of this process. FDA tools allow us to summarize and capture variability within this process. To achieve this, we perform a basis decomposition using multivariate functional principal components analysis (MFPCA), and then describe players’ shooting styles based on this decomposition. Finally, representing player data onto the basis decomposition allows us to use clustering techniques to identify players with similar shooting behaviors. This approach to analyzing shooting patterns provides nuanced insights that can be valuable for coaches, analysts, and players alike.
@@ 2 Dataset
Our motivating dataset is derived from the NBA. To access comprehensive game data, we use the Python package nba_api [^1] https://github.com/swar/nba_api/, which interfaces with the APIs provided by nba.com. The dataset includes both made and missed field goal attempt locations from the offensive half-court, covering all NBA games from the $2018-2019$ to the $2023-2024$ regular seasons. The offensive half-court is a $50$ ft (sideline to sideline) by $47$ ft (baseline to mid court line) rectangle. Filtering the dataset, we focus on players who attempted more than $1000$ field goals during this six-season period, resulting in a cohort of $173$ players out of the $573$ players who have stepped onto the court during these seven years. Collectively, these players attempted $717381$ shots, of which $340466$ ($47\%$) were successful. To ensure data quality, we exclude attempts classified as impossible (e.g., out-of-bounds), leaving a dataset of $716114$ shots, with $340430$ ($48\%$) successful attempts.
In the NBA, players are typically categorized by their roles on the court into three primary positions: ‘guard’, ‘forward’, and ‘center’. Guards are responsible for ball handling, playmaking, and perimeter shooting. Centers, often referred to as ”big men”, primarily operate near the basket to score in the paint. Forwards are versatile players who contribute in multiple areas of the court. Additionally, the NBA recognizes hybrid roles for players with overlapping skill sets: ‘guard-forward’, ‘forward-guard’, ‘center-forward’, and ‘forward-center’. To ensure a sufficient number of players in each group, we aggregate the ‘guard-forward’ and ‘forward-guard’ roles into a single ‘forward-guard’ category, and the ‘center-forward’ and ‘forward-center’ roles into a single ‘forward-center’ category. This results in five distinct positional groups (‘guard’, ‘forward-guard’, ‘forward’, ‘forward-center’ and ‘center’), corresponding to the number of players on the court. The number of players across these categories is shown in Table 1.
@@ 3 Methodology
The first step in our approach is to convert the raw shots charts into smooth functions suitable for analysis. To achieve this, we use a two-dimensional kernel density estimator with a Gaussian kernel for both missed and made shots. After normalizing the court to the $[0,1]\times[0,1]$ range, bandwidth estimation is performed utilizing Silverman’s rule (Silverman,, 1986). The density estimation is conducted on a regularly sampled grid of $201\times 201$ points. We then represent the smoothed densities of missed and made shots for each player as the two features of a bivariate functional observation. Both features are defined on the same two-dimensional rectangular domain and observed over the $201\times 201$ grid. Note that the choice of the grid resolution is arbitrary and can be adjusted as needed. Figure 2 shows the density estimates of the same five players whose shots charts are presented in Figure 1.
We apply multivariate functional principal components analysis (Happ and Greven,, 2018) to decompose each estimated density into a common set of basis functions and player-specific coefficients. Let $X$ denote an observation of the densities of a player, represented as a vector of two functions: the first corresponding to the density of missed shots and the second to the density of the made shots. We aim to find basis functions $\{\phi_{k}\}_{1\leq k\leq K}$ that provide the most parsimonious representation of $X$, for a given $K$, as
@@
where $\mu(s,t)$ is the mean (average) function, and $c_{k}$ are the player-specific basis coefficients, referred to as scores. Each eigenfunction in the decomposition (1) characterize a variation away from the mean function, with the player-specific scores indicating the magnitude and direction of the deviation for each individual player. We estimate four components ($K=4$), based on the approach of Golovkine et al., (2024), explaining approximately $90\%$ of the total variance in the data. The scores $c_{k}$ are estimated as the projection of the observations $X$ onto the eigenfunctions $\phi_{k}$. These scores provide a finite-dimensional representation of the functional data, capturing player-specific characteristics, and can be used as scalar variables in subsequent analyses. To create a data-driven taxonomy of players based on MFPCA, we apply a $k$-medoids clustering algorithm (Kaufman and Rousseeuw,, 2009) to the normalized functional principal component scores $c_{k}$, using $k=5$ clusters, as the number of players on the court. The $k$-medoids algorithm minimizes the dissimilarity between each player’s scores and the medoid of its cluster. We use the weighted euclidean distance as the measure of dissimilarity. This method offers a practical advantage over the $k$-means clustering algorithm: the center of each cluster is an actual player from the dataset rather than an averaged representation of all players within the cluster, enhancing interpretability. In our analysis, we consider two weighting schemes to calculate the distance: one assigns equal weights to all components, and the other weights components given the proportion of variance they explain.
Figure 2: Smoothed shot chart densities for five selected NBA players. The densities have been individually normalized between 00 and $1$.
@@ 4.1 MFPCA decomposition of the densities
We apply MFPCA to estimate the mean functions and eigenfunctions, providing a description of players shooting behavior. Figure 3 presents the estimation of the mean functions alongside the estimated functional principal components, which capture deviations from the mean shooting density. Unlike the shooting densities, the functional principal components may take on negative values as they represent deviations from the mean rather than densities themselves. For presentation purposes, the functional principal components have been normalized between $-1$ and $1$. Blue areas correspond to negative values of the functional components, and thus contribute negatively to the density, while red area corresponds to positive values of the functional components, and thus contribute positively to the density. The mean chart represents the shooting behavior of the average player in the NBA. Shots are most frequently taken in the paint, then the most common position is behind the three-point line, in the wings and the corners. The elbows and the short-corners are less important as shooting spots. The estimated functional components can be explained as characterizing different shooting behaviors. We first note that, except for $\phi_{3}$, the functional components for missed and made shots are similar. This suggests that players do not have specific shooting position where the percentage of missed shots significantly exceeds that of made shots. The first component, explained $67.9\%$ of the total variance, contrasts players based on their overall shooting frequency relative to the average NBA player. A player with a positive score $c_{1}$ for $\phi_{1}$ will shot more frequently, both successfully and unsuccessfully, than average. These effects are concentrated the high-density areas of the court such as the paint and beyond the arc (i.e., $\phi_{1}$ has a similar shape to the mean function). Conversely, a negative score corresponds to players who shoot less frequently than the average, with effects concentrated in these same regions. The second component (explaining $8.5\%$ of the total variance) highlights a contrast between shooting inside and outside the three-pointer line. Players with a positive score $c_{2}$ for $\phi_{2}$ prefer to shoot closer to the basket, while those with negative scores prefer long-range shots. The third component (explaining $6.1\%$ of the total variance) contrasts players that shoot in the paint and the others. Positive scores $c_{3}$ are associated with frequent attempts and successes in the paint, while negative scores would prefer to shoot farther out. Interestingly, players who perform well near the basket also tend to attempt and miss shots from the corners. The fourth component (explaining $4.0\%$ of the total variance) captures differences between shooting from the wings and shooting along the axis of the hoop. Players with positive (resp. negative) score will shot in the axis of the hoop (resp. from the wing).
@@
Figure 3: Estimates of the mean surface and the eigenfunctions of the shots density charts. The eigenfunctions represent deviations from the mean surface and are defined up to a sign, making their values at specific points not directly interpretable as densities. Note that the mean surface and eigenfunctions have been normalized to be between $-1$ and $1$ for presentation purposes.
@@
As an example, Figure 4 illustrates the decomposition of Stephen Curry’s density shot chart. Curry shoots more frequently than the average player, as indicated by his positive coefficient $c_{1}$, reflecting a higher number of both missed and made shots compared to the average player. His second coefficient $c_{2}$ is negative and close to zero, suggesting only a slight preference for shooting beyond the three-point line. The small negative value of his third coefficient $c_{3}$ indicates shooting behavior close to the average for this component. However, Curry prefers to shoots in the axis of the basket (fourth coefficient $c_{4}$ is positive) compared to the other players.
@@
Figure 5 presents the estimated scores $c_{k}$ from (1) colored by the NBA-defined playing positions. One representative player from each position is highlighted (same as Figures 1 and 4). Notably, the scores do not separate players by position. From Figure 5, it can be seen that some players exhibit scores (coefficients) that substantially differ from the general distribution. For example, for the second eigenfunction, Chris Paul (with a large coefficient) and Duncan Robinson (with a large negative coefficient) may be designed as outliers. Similarly, for the third eigenfunction, Chris Paul has a large negative coefficient. Although a positive $c_{2}$ typically indicates a preference for shooting near the basket and a negative $c_{3}$ suggests a preference for shooting beyond the three-point line, Chris Paul’s scoring pattern appears contradictory. This inconsistency arises because Chris Paul’s most common shooting position is the high post, located between the three-point line and the painted area. In this case, the eigenfunctions compensate for one another to capture his shooting style, which is relatively rare in the modern NBA. Finally, for the fourth eigenfunction, Reggie Bullock Jr. has a large negative coefficient, indicating a strong preference for shooting from the wings and corners.
Figure 5: Scatter plots of the player-specific basis coefficients (scores) colored using the NBA defined typology of players. We highlighted one player per position. The scores have been jittered on the y-axis for visualization purposes.

### v2

# Is Stephen Curry really a guard? A new perspective on player typologies using functional data analysis

@@ 1 Introduction
With the increasing availability of large, detailed, and complex sports performance datasets, sports analytics has become a rapidly expanding field of both research and application, drawing significant interest from the statistics and machine learning communities. In team sports, traditional approaches have often focused on predicting game outcomes based on team characteristics (dixonModellingAssociationFootball1997; babootaPredictiveAnalysisModelling2019); for recent reviews, see horvatUseMachineLearning2020; bunkerApplicationMachineLearning2022. More recently, the availability of tracking data has enabled researchers to delve into more specific aspects of various sports, such as evaluating individual player performance in soccer (pappalardoPlayeRankDatadrivenPerformance2019; decroosActionsSpeakLouder2019) or analyzing defensive performance in American football passing plays (schmidSimulatingDefensiveTrajectories2021). See kovalchikPlayerTrackingData2023 for a recent review on machine learning approaches to analyze tracking data. Basketball, in particular, presents opportunities for analysis, supported by extensive data provided by the National Basketball Association (NBA).
@@
Basketball is a team sport in which two teams of five players compete to score points by shooting a ball through the opponent’s hoop while defending their own. The game has been extensively studied through various analytical perspectives. For example, researchers have examined shooting mechanics and movements (millerRelationshipBasketballShooting1996; liuChangesBasketballShooting1999; okazakiReviewBasketballJump2015), estimated pre-game and in-game winning probabilities (maddoxBayesianEstimationIngame2022), and modeled outcomes of possessions (cervoneMultiresolutionStochasticProcess2016). Statistical models have also been applied to analyze defensive strategies (csataljayEffectsDefensivePressure2013) and spatial performance on the court (zuccolottoSpatialPerformanceAnalysis2023). For a comprehensive review of performance analysis in basketball, see ternerModelingPlayerTeam2021.
@@
In basketball, game outcomes are heavily influenced by shooting accuracy, as a player’s effectiveness is often defined by their ability to consistently score from various court locations. Analyzing shooting patterns thus serves as a powerful tool, not only for assessing individual performance but also for refining team strategies. An important tool for examining shooting patterns is shooting charts, which are visual representations that display all the shots a player has attempted and made over a season or a single game. Shooting charts are extensively studied in basketball research, with much of the work focusing on modeling them as spatial point processes. This approach typically involves creating a count matrix of shots taken by each player on a discretized court, then estimating an intensity surface over this grid by fitting a log-Gaussian Cox process. A low-rank approximation of these intensity surfaces is achieved using non-negative matrix factorization, as described in millerFactorizedPointProcess2014; franksCharacterizingSpatialStructure2015. Non-parametric estimation of the intensity surface is also considered in yinBayesianNonparametricLearning2022. These low-rank intensity surfaces help characterize NBA players’ shooting tendencies and evaluate their scoring efficiency. Some models do not require the intensity surface fitting approach. For example, reichSpatialAnalysisBasketball2006 employed conditionally autoregressive models to capture spatial correlations, while huZeroinflatedPoissonModel2023 fitted a Bayesian zero-inflated Poisson model to represent shooting attempts for players with distinct shooting styles. Although these methods effectively summarize player data, they are less suitable for comparing different players.
@@
To compare players’ shooting behaviors, specialized methods have been developed. fuHoopInSightAnalyzingComparing2024 employ (interactive) data visualization techniques to analyze and contrast shooting performance. In a different approach, huBayesianGroupLearning2021 use a log-Gaussian Cox process to model the shooting intensity with a specific similarity measure for comparing players. jiaoBayesianMarkedSpatial2021 propose a Bayesian marked spatial point process model to capture shot intensity, investigating whether players’ success rates are higher in areas where they attempt more shots and clustering players based on this pattern. In another study, wong-toiJointAnalysisField2022 analyze shooting patterns across players by dividing the court into a grid and using a Bayesian nonparametric mixture model to identify latent clusters among players. Similarly, yinAnalysisProfessionalBasketball2023 apply a Bayesian nonparametric matrix clustering technique to estimate intensity maps, offering insights into shot selection characteristics. All these methods require a discretized representation of the court.
@@
To overcome the requirement of court discretization, we propose to estimate the shot density continuously across the offensive half-court and model these densities as functional data (ramsayFunctionalDataAnalysis2005; kokoszkaIntroductionFunctionalData2017). This idea first appeared in our paper golovkineUseGramMatrix2024. Both missed and made shot charts are treated as multivariate functional data over a two-dimensional domain (i.e., the offensive half-court). Functional Data Analysis (FDA) addresses the analysis of realizations of a continuous stochastic process, observed at discrete sampling points and defined on compact domains. Densities are computed separately for each shot type (made and missed) and each player and are therefore normalized for each individuals total number of each shots. We prefer this over an un-normalized representation that contains information about overall shot volume, so that we can make make fair comparisons of shooting patterns that are not affected by, e.g., number of games or minutes played. By treating shot density as a stochastic process defined over the offensive half-court, we can consider each player’s shot density as a realization of this process. FDA tools allow us to summarize and capture variability within this process. To achieve this, we perform a basis decomposition using multivariate functional principal components analysis (MFPCA), and then describe players’ shooting styles based on this decomposition. Finally, representing player data onto the basis decomposition allows us to use clustering techniques to identify players with similar shooting behaviors. This approach to analyzing shooting patterns provides nuanced insights that can be valuable for coaches, analysts, and players alike.
@@ 2 Dataset
Our motivating dataset is derived from the NBA. To access comprehensive game data, we use the Python package nba_api [^1] https://github.com/swar/nba_api/, which interfaces with the APIs provided by nba.com. The dataset includes both made and missed field goal attempt locations from the offensive half-court, covering all NBA games from the 201820192018-2019 to the 202320242023-2024 regular seasons. The offensive half-court is a $50$ ft (sideline to sideline) by $47$ ft (baseline to mid court line) rectangle. Filtering the dataset, we focus on players who attempted more than $1000$ field goals during this six-season period, resulting in a cohort of $173$ players out of the $573$ players who have stepped onto the court during these six seasons. Collectively, these players attempted $717381$ shots, of which $340466$ ($47\%$) were successful. To ensure data quality, we exclude attempts classified as impossible (e.g., out-of-bounds), leaving a dataset of $716114$ shots, with $340430$ ($48\%$) successful attempts.
In the NBA, players are typically categorized by their roles on the court into three primary positions: ‘guard’, ‘forward’, and ‘center’. Guards are responsible for ball handling, playmaking, and perimeter shooting. Centers, often referred to as “big men”, primarily operate near the basket to score in the paint. Forwards are versatile players who contribute in multiple areas of the court. Additionally, the NBA recognizes hybrid roles for players with overlapping skill sets: ‘guard-forward’, ‘forward-guard’, ‘center-forward’, and ‘forward-center’. To ensure a sufficient number of players in each group, we aggregate the ‘guard-forward’ and ‘forward-guard’ roles into a single ‘forward-guard’ category, and the ‘center-forward’ and ‘forward-center’ roles into a single ‘forward-center’ category. This results in five distinct positional groups (‘guard’, ‘forward-guard’, ‘forward’, ‘forward-center’ and ‘center’), corresponding to the number of players on the court. The number of players across these categories is shown in Table 1.
@@ 3 Methodology
The first step in our approach is to convert the raw shot charts into smooth functions suitable for analysis. To achieve this, we use a two-dimensional kernel density estimator with a Gaussian kernel for both missed and made shots. After normalizing the court to the $0,1]\times[0,1]$ range, bandwidth estimation is performed utilizing Silverman’s rule (silvermanDensityEstimationStatistics1986) for each component of each player. The density estimation is conducted on a regularly sampled grid of $201\times 201$ points. We then represent the smoothed densities of missed and made shots for each player as the two features of a bivariate functional observation. Both features are defined on the same two-dimensional rectangular domain and observed over the $201\times 201$ grid. Note that the choice of the grid resolution is arbitrary and can be adjusted as needed. Figure 2 shows the density estimates of the same five players whose shot charts are presented in Figure 1.
After the density estimation, our data consist of independent trajectories of a vector-valued stochastic process $X=(X^{(1)},X^{(2)})^{\top}$. Let $\mathcal{T}=[0,1]\times[0,1]$. The first coordinate $X^{(1)}$ corresponds to the density of missed shots and the second coordinate $X^{(2)}$ corresponds to the density of made shots. Both of them are assumed to belong to $\mathcal{L}^{2}(\mathcal{T})$, the Hilbert space of square-integrable real-valued functions defined on $\mathcal{T}$ having the usual inner product. Thus, $X$ is a stochastic process indexed by $\mathbf{t}=(t_{1},t_{2})$ belonging to the $2$-fold Cartesian product $\mathcal{T}\times\mathcal{T}$ and taking values in the $2$-fold Cartesian product space $\mathcal{H}=\mathcal{L}^{2}(\mathcal{T})\times\mathcal{L}^{2}(\mathcal{T})$. In this space, we define the inner product as the sum of the inner products in $\mathcal{L}^{2}(\mathcal{T})$. Let $\mu:\mathcal{T}\times\mathcal{T}\rightarrow\mathcal{H}$ denote the mean function of the process $X$ and let $C$ denote the matrix-valued covariance function of the process $X$. We can then define the covariance operator of $X$, $\Gamma:\mathcal{H}\rightarrow\mathcal{H}$, as an integral operator with kernel $C$. Assuming that the covariance operator $\Gamma$ is a compact positive operator on $\mathcal{H}$ and using the results in happMultivariateFunctionalPrincipal2018, there exists a complete orthonormal basis, the eigenfunctions, $\{\phi_{k}\}_{k\geq 1}\subset\mathcal{H}$ associated to a set of real numbers, the eigenvalues, $\{\lambda_{k}\}_{k\geq 1}$ that satisfy
$\Gamma\phi_{k}=\lambda_{k}\phi_{k},\quad\text{and}\quad\lambda_{k}\rightarrow 0\quad\text{as}\quad k\rightarrow\infty.$
Using the multivariate Karhunen-Loève theorem (happMultivariateFunctionalPrincipal2018), we obtain the decomposition
$X(s,t)=\mu(s,t)+\sum_{k=1}^{\infty}c_{k}\phi_{k}(s,t),\quad s,t\in[0,1]\times[0,1],$ (1)
where $c_{k}$ are the projections of the centered curves onto the eigenfunctions.
We apply multivariate functional principal components analysis (happMultivariateFunctionalPrincipal2018) to decompose each estimated density into a common set of basis functions and player-specific coefficients using (1). We aim to find basis functions $\{\phi_{k}\}_{1\leq k\leq K}$ that provide the most parsimonious representation of $X$, for a given $K$, as
@@
where $\mu(s,t)$ is the mean (average) function, and $c_{k}$ are the player-specific basis coefficients, referred to as scores. Each eigenfunction in the decomposition (2) characterizes a variation away from the mean function, with the player-specific scores indicating the magnitude and direction of the deviation for each individual player. We estimate four components ($K=4$), based on the approach of golovkineUseGramMatrix2024, explaining approximately $90\%$ of the total variance in the data. The scores $c_{k}$ are estimated as the projection of the observations $X$ onto the eigenfunctions $\phi_{k}$ using the inner product in $\mathcal{H}$. These scores provide a finite-dimensional representation of the functional data, capturing player-specific characteristics, and can be used as scalar variables in subsequent analyses. To create a data-driven taxonomy of players based on MFPCA, we apply a $k$-medoids clustering algorithm (kaufmanFindingGroupsData2009) to the normalized functional principal component scores $c_{k}$, using $k=5$ clusters, as the number of players on the court. Moreover, we will compare the estimated clusters to the $k=5$ nominal positions provided by the NBA. We do not expect perfect agreement – defence, play-making and movement all play a role in addition to shooting. However, because shooting behavior is a key component of a player’s overall game, the positions comparison can help inform the labels we provide to our clusters. The $k$-medoids algorithm minimizes the dissimilarity between each player’s scores and the medoid of its cluster. We use the weighted Euclidean distance as the measure of dissimilarity. This method offers a practical advantage over the $k$-means clustering algorithm: the center of each cluster is an actual player from the dataset rather than an averaged representation of all players within the cluster, enhancing interpretability. In our analysis, we consider two weighting schemes to calculate the distance: one assigns equal weights to all components, and the other weights components given the proportion of variance they explain.
Figure 2: Smoothed shot chart densities for five selected NBA players. The densities have been individually normalized between 0 and $1$.
@@ 4.1 MFPCA decomposition of the densities
We apply MFPCA to estimate the mean functions and eigenfunctions, providing a description of players’ shooting behavior. Figure 3 presents estimates of the mean functions alongside the estimated functional principal components, which capture deviations from the mean shooting density. Unlike the shooting densities, the functional principal components may take on negative values as they represent deviations from the mean rather than densities themselves. For presentation purposes, the functional principal components have been normalized between 1-1 and $1$. Blue areas correspond to negative values of the functional components, and represent negative deviations from the mean density, while red areas correspond to positive values of the functional components, and thus positive deviations. Because the individual densities have been normalized separately for each outcome (made and missed) and player, the patterns captured by the mean and functional principal components should be interpreted as reflecting where players tend to take made and missed shots, not where shots are more or less likely to be made (i.e., success probability or efficiency).
The mean chart represents the average shooting behavior of NBA players and is displayed in the first column of Figure 3. For shots made (second row, first column), the mean density displays very high values in the small region under the paint. This indicates that a large proportion of the shots that players make successfully are taken from this area. The next highest density region in the mean shots made chart is around the perimeter of the three-point line, but this density is lower than under the paint. For shots missed (first row, first column), there is also a high density region under the paint, indicating that a large number of the missed shots are also taken from this area. However, in comparison with the shots made density, the next highest density region – around the perimeter of the three-point line – is larger and closer in density values to the area under the paint. Therefore, we can deduce that although a large number of both made shots and missed shots are taken from under the paint and around the three point area, a comparatively larger proportion of missed shots are taken from the three point line. This difference is intuitive, because three-point shots are inherently harder to make. Overall, the elbows and the short-corners are less important as shooting spots because these regions account for a smaller number of both missed and made shots.
@@
The estimated functional components can be understood as characterizing different shooting behaviors. We first note that, except for $\phi_{3}$, the functional components for missed and made shots are similar. This suggests that the dominant patterns of player-to-player variation are driven by shot location, regardless of whether the shots are made or missed. The first component explains $67.9\%$ of the total variance and is similar in shape to the mean densities because it assigns higher positive density values to the areas under the paint and around the outside of the perimeter arc. Although the average player tends to have a large proportion of their made and missed shots from this area, players with a positive score $c_{1}$ tend to make and miss even more of their shots than average from these regions. In contrast, players with a negative score $c_{1}$ have a smaller-than-average proportion of their made and missed shots taken from these “typical” shooting regions. In addition, for the missed shots, there is a small, compact blue area, appearing as a small dot, in the location directly under the basket. This means that players with a positive score $c_{1}$ have fewer of their missed shots from this location. The second component, explaining $8.5\%$ of the total variance, highlights a contrast between shooting inside and outside the three-point line. Players with a positive score $c_{2}$ for $\phi_{2}$ prefer to shoot closer to the basket, while those with negative scores prefer long-range shots. The third component (explaining $6.1\%$ of the total variance) contrasts players that shoot in the paint and the others, while also capturing some variation in missed shots from the corners. For made shots, positive scores $c_{3}$ are associated with frequent attempts and successes in the paint and fewer made shots from further out. For missed shots, positive scores $c_{3}$ also indicate frequent attempts and misses from around the paint, but interestingly also capture high proportions of missed shots from the corners, indicated by the two red regions on either side. This pattern indicates that these players who tend to take a large number of their shots (both made and missed) from inside the paint also tend to take shots from the corners but are generally unsuccessful (i.e., we see the high densities in the corners for the missed shots but not the made shots). Finally, the fourth component accounts for $4.0\%$ of the total variance and characterizes differences in shooting centrally (in the paint and along the central portion of the three-point arc) and shooting from the wings. Players with a positive (negative) score $c_{4}$ tend to make and miss more of their shots from central regions of the court, and fewer of their shots from the wings.
@@
Figure 3: Estimates of the mean surface and the eigenfunctions of the shots density charts. The eigenfunctions represent deviations from the mean surface and are defined up to a sign, making their values at specific points not directly interpretable as densities. Note that the mean surface and eigenfunctions have been normalized to be between 1-1 and 11 for presentation purposes.
@@
As an example, Figure 4 illustrates the decomposition of Stephen Curry’s density shot chart. Curry shoots more frequently than the average player, as indicated by his positive coefficient $c_{1}$, reflecting a higher number of both missed and made shots compared to the average player. His second coefficient $c_{2}$ is negative and close to zero, suggesting only a slight preference for shooting beyond the three-point line. The small negative value of his third coefficient $c_{3}$ indicates shooting behavior close to the average for this component. However, Curry prefers to shoot in the axis of the basket (fourth coefficient $c_{4}$ is positive) compared to the other players.
@@
Figure 5 presents the estimated scores $c_{k}$ from (2) colored by the NBA-defined playing positions. One representative player from each position is highlighted (same as Figures 1 and 4). Notably, the scores do not separate players by position. From Figure 5, it can be seen that some players exhibit scores (coefficients) that substantially differ from the general distribution. For example, for the second eigenfunction, Chris Paul (with a large coefficient) and Duncan Robinson (with a large negative coefficient) may be designated as outliers. Similarly, for the third eigenfunction, Chris Paul has a large negative coefficient. Although a positive $c_{2}$ typically indicates a preference for shooting near the basket and a negative $c_{3}$ suggests a preference for shooting beyond the three-point line, Chris Paul’s scoring pattern appears contradictory. This inconsistency arises because Chris Paul’s most common shooting position is the high post, located between the three-point line and the painted area. In this case, the eigenfunctions compensate for one another to capture his shooting style, which is relatively rare in the modern NBA. Finally, for the fourth eigenfunction, Reggie Bullock Jr. has a large negative coefficient, indicating a strong preference for shooting from the wings and corners.
Figure 5: Scatter plots of the player-specific basis coefficients (scores) colored using the NBA defined typology of players. We highlighted one player per position. The scores have been standardized and jittered on the y-axis for visualization purposes. The first component discriminates players by their overall shooting volume, with positive scores indicating a higher-than-average number of shots. The second component contrasts between two-pointers and three-pointers, with positive scores associated with shots near the basket. The third component differentiates players that shots in the paint, combined with frequent missed shots from the corners and associated to positive scores, and the others. The fourth component contrasts between centrally located shooting and wing-area shooting, with positive scores indicating more central shooting.